# Quantum-Enhanced Processing with Tensor-Network Frontends for Privacy-Aware Federated Medical Diagnosis

### 02 Apr 2026 v1

@@ IV-A Experiment 1: Quantum Enhancement across MPS, TTN, and MERA
The first experiment characterizes how the Quantum-Enhanced Processor (QEP) behaves across different client-side tensor-network architectures. Specifically, we compare Classical and Quantum modes for MPS, TTN, and MERA in order to determine whether the contribution of the QEP depends on the latent structure induced by the tensor-network frontend.
In the Classical mode, the server-side feature after gating is passed directly to the classifier. In the Quantum mode, the same latent feature is processed by the QEP, which generates observable-based quantum statistics, classically decodes them into the shared latent space, and adaptively fuses them with the original classical representation. This comparison isolates the representational contribution of the quantum-enhanced branch under otherwise matched training conditions.
We evaluate both inference performance and internal QEP behavior. Inference is reported under both the standard decision threshold and a validation-optimized threshold derived from Youden’s $J$ statistic and applied to the test split. Internal analysis tracks the mean fusion coefficient $\alpha$, the dispersion of the quantum-enhanced branch through $q$-std, and their joint behavior across training. This design allows us to assess not only whether the QEP changes output metrics, but also how its internal operating regime depends on the tensor-network frontend.
To further characterize the behavior of the QEP in the most favorable frontend setting observed in Experiment 1, we additionally analyze the TTN-based model with respect to qubit scaling and noise sensitivity. For qubit scaling, the QEP is instantiated with $N_{q}\in\{4,6,8,10,12,14,16\}$, and the resulting test-accuracy distributions are compared against the classical TTN baseline. For noise sensitivity, we fix $N_{q}=8$ and compare the classical TTN baseline with four quantum execution conditions: noiseless, depolarizing, thermal, and mixed noise. Both analyses are reported as distributional summaries of test accuracy over repeated runs, allowing comparison of both central tendency and variability across conditions.
@@ IV-B Experiment 2: MPC Mode Benchmark
The second experiment benchmarks the communication cost of MPC-secured aggregation under progressively stronger protection settings. The benchmark isolates the server-side communication induced by protected latent aggregation and is therefore not intended as a full accounting of all end-to-end distributed traffic. Rather, it provides a controlled analysis of how protected latent dimension and security setting determine the communication burden of the secure aggregation stage itself.
We vary the number of clients from $n=1$ to $30$. For each value of $n$, we consider seven scenarios. Scenario 0 is an insecure baseline in which a central server $S$ performs the post-client aggregation and transformation directly. Scenarios 1–3 correspond to passive-security MPC settings with progressively richer protected functionality: Scenario 1 protects only weighted aggregation, Scenario 2 additionally protects normalization, and Scenario 3 further protects the post-aggregation transformation stage. Scenarios 4–6 are the corresponding active-security counterparts. Table I summarizes the security setting and protected functionality associated with each scenario.
@@ IV-B1 Cost Model and Measurement Protocol
We employ a symbolic cost meter to quantify the communication cost of each scenario. The meter tracks three types of communication: (i) data sent by clients to computation nodes, (ii) data exchanged between computation nodes, and (iii) data transmitted during output reconstruction.
In all scenarios, we assume $k=64$, corresponding to an encoding in which client inputs fit into 64-bit machine words. For Scenarios 2 and 5, division is metered using $\theta=5$. This benchmark is intended to capture relative scaling trends across MPC scenarios rather than protocol-tight cryptographic cost bounds. Additional implementation overheads that may arise in practice are not included in the present cost model.
@@ IV-B2 Fairness and Comparability Across Modes
To avoid conflating security modeling with numerical instability, we assume that fixed-point parameters can be chosen so that the accuracy of the MPC computation matches that of the original insecure floating-point computation. As noted in [26], automated tools can assist in fixed-point parameter selection. Under this assumption, the benchmark should be interpreted as a comparison of secure-computation overheads under matched numerical behavior, rather than as a study of accuracy degradation caused by quantization or approximation artifacts.
@@ V-A Results of Experiment 1: Quantum Enhancement across MPS, TTN, and MERA
Figure 4: Threshold-optimized test performance of the Classical and Quantum modes across MPS, TTN, and MERA. Boxplots report class-wise Precision, Recall, F1-score, and overall Accuracy.Results are shown for the default QEP setting ($N_{q}=16$).
@@
Figure 6: TTN-based Quantum-Enhanced Processor (QEP): (a) test-accuracy distributions under qubit-count scaling ($N_{q}=4$ to 16), and (b) robustness under representative noise conditions at $N_{q}=8$. The classical TTN baseline is included for reference.
@@
Fig. 4 and 5 summarize the inference performance and internal QEP behavior in the comparison between Classical and Quantum modes across MPS, TTN, and MERA. Overall, the results indicate that the effect of quantum enhancement is architecture-dependent rather than uniform across tensor-network frontends.
Under threshold-optimized evaluation, the combination of tensor-network architecture and post-aggregation quantum enhancement clearly affects the performance profile. The main effect of the QEP is not a uniform increase in overall Accuracy, but a redistribution of class-wise Precision, Recall, and F1 across frontend types. In the present study, TTN exhibits the most balanced overall profile, achieving strong Accuracy and Pneumonia F1 while remaining competitive on the remaining class-wise metrics. MPS and MERA also remain competitive, but their gains under quantum enhancement are less consistently aligned with the strongest downstream performance.
From a clinical perspective, an important observation is that Pneumonia Recall remains high across all architectures, at around 0.90. This suggests that the proposed hybrid pipeline maintains strong sensitivity to positive cases even when the frontend architecture changes. In screening-oriented settings, where missed pneumonia cases are often more critical than moderate changes in aggregate balanced metrics, this is an encouraging property of the overall framework.
The internal diagnostics further show that the QEP operates differently depending on the frontend. Although the mean fusion coefficient $\alpha$ decreases over training for all three architectures, the resulting operating regimes differ substantially. MPS reaches the highest final $q$-std, MERA remains intermediate, and TTN stabilizes at a comparatively lower and more controlled quantum-branch dispersion. This indicates that the effect of the QEP is not explained by quantum-branch variability alone. Instead, the results suggest that learning performance is shaped by the interaction between frontend topology and post-aggregation quantum refinement, with TTN providing the most favorable balance in the present setting.
Fig. 6 summarizes additional analyses for the TTN-based QEP, focusing on qubit-count scaling and noise robustness. Because TTN provides the most favorable overall profile in the frontend comparison, we use it here as the reference setting for a more detailed examination of the operating range of the quantum branch.
@@
The left panel shows the test-accuracy distributions of the TTN+QEP model for qubit counts from 4 to 16, together with the classical TTN baseline. The lower-qubit settings, especially $N_{q}=4$ and $N_{q}=6$, tend to show lower or less favorable accuracy distributions than the higher-qubit settings. By contrast, from $N_{q}\geq 8$, the distributions appear more stable and remain within a broadly similar range. In the present TTN-based setting, this suggests that the QEP operates more reliably once the qubit count is sufficiently matched to the 64-dimensional latent representation.
@@
The right panel shows the test-accuracy distributions for the TTN+QEP model at $N_{q}=8$ under different noise conditions. Relative to the noiseless setting, all noisy conditions exhibit some degree of performance degradation, although the extent of degradation varies across noise types. The depolarizing setting remains comparatively close to the noiseless case, whereas the thermal setting shows a wider spread and includes lower-performing runs, indicating stronger sensitivity under that noise model. The mixed-noise setting also shows a degradation trend relative to the noiseless setting while retaining a broadly comparable overall range. Taken together, these results suggest that the TTN-based QEP is affected by noise in a noise-type-dependent manner.
@@ V-B Results of Experiment 2: MPC Benchmark
Fig. 7(a) and 7(b) summarize the modeled communication overhead for the passive and active MPC scenarios, together with the insecure baseline.
@@
The results show that, under the present abstraction of server-side MPC communication, the dominant communication cost is governed by the bit-width $k$, the protected representation dimension, and the selected protection scenario, rather than by the number of aggregated clients itself. Because the benchmark isolates the communication induced by MPC processing of protected latent representations, client count does not appear as the dominant scaling factor in the present formulation. In both figures, $k=64$ is fixed while the protected representation dimension varies from the raw dimension $D_{\mathrm{raw}}=784$ to the compressed latent dimension $d=64$.
Active versus passive security.
Active security preserves the same qualitative scaling trend as passive security while introducing a moderate multiplicative overhead. In the present protocol model, this additional cost is limited, although different protocol assumptions could lead to substantially larger overheads.
Effect of representation compression.
The dominant communication trend is governed by the representation dimension. This indicates that client-side compression has a direct impact on the effective cost of secure aggregation: reducing the latent dimension before MPC reduces the communication surface on which security overhead is applied.
Varying the number of clients.
Because the present benchmark isolates the communication cost of the MPC stage itself, the number of clients does not emerge as the dominant scaling factor in the reported curves. This is also true for the insecure baseline when only the output-side aggregation communication is considered. Even if client-to-aggregator transmission were explicitly included, the main conclusion would remain unchanged: under the linear secret-sharing model considered here, the dominant cost trend is governed primarily by the protected representation dimension and the protection scenario rather than by the number of aggregated clients.
@@ VI Discussion
From a clinical perspective, the consistently high Pneumonia Recall across architectures is encouraging. In screening or triage settings, sensitivity to pneumonia-positive cases is often more important than small fluctuations in class-balanced aggregate metrics, because false negatives carry disproportionate clinical cost. The present results therefore suggest that hybrid TN+QEP pipelines can preserve clinically relevant sensitivity while still allowing architectural differences to shape the overall operating point.
@@ VII Conclusion
In this work, we investigated the role of the QEP within a privacy-aware federated learning framework for medical image classification. The proposed architecture combines client-side tensor-network representation learning, an abstract secure-aggregation model for MPC benchmarking, and a post-aggregation quantum refinement module based on observable expectation values, classical decoding, differentiable bypass fusion, and adaptive residual mixing.
@@
The additional qubit-scaling and noise analyses further clarify the practical behavior of the QEP. For the present 64-dimensional latent input, performance becomes more stable from $8$ qubits onward, while noisy conditions at $8$ qubits degrade performance relative to the noiseless case in a noise-type-dependent manner. These results do not establish hardware-level benefit, but they help characterize the operating range and robustness of the hybrid quantum branch within the proposed pipeline.
@@
While the present results are obtained in a simulator-based regime, an important motivation of the proposed framework lies in future hardware-oriented use beyond the classically simulable range. When quantum processors with on the order of tens of usable qubits are employed in regimes where full classical simulation is no longer practical, it becomes meaningful to apply quantum processing to tensor-network-compressed latent features rather than to the original large-scale input. In this sense, compressing very high-dimensional data—for example, from hundreds of thousands of dimensions to a few hundred latent features—may provide a realistic pathway for practical quantum-enhanced learning on future devices.
Future work will proceed in three directions. First, implementing end-to-end cryptographic MPC with explicit fixed-point execution will allow direct validation of the present cost model and a more detailed characterization of secure learning performance. Second, protocol-level optimization will be necessary to improve scalability and reduce the overhead of stronger security scenarios. Third, further study of the QEP—including deeper circuits, stronger ablation against capacity-matched classical surrogates, and more realistic noisy or hardware-backed settings—will be important for clarifying when hybrid quantum enhancement provides meaningful benefit in privacy-aware federated learning.
@@ Acknowledgments
The authors acknowledge helpful conversations with Mark Medum Bundgaad, Miyoji Kakinuki, Akiko Kamigori, Kazufumi Okazaki, Tomah Sogabe, Naoki Yamamoto, Eriko Kaminishi, Toshiki Yasuda. Hiroshi Yamauchi acknowledges the support of Yoshi-aki Shimada, Yosuke Komiyama and Ryuji Wakikawa. This work was partly supported by NEDO Challenge Quantum Computing “Solve Social Issues!”.

### 03 Apr 2026 v2

@@ IV-A Experiment 1: Quantum Enhancement across MPS, TTN, and MERA
The first experiment examines whether the effect of the Quantum-Enhanced Processor (QEP) depends on the client-side tensor-network frontend. We compare Classical and Quantum modes for MPS, TTN, and MERA under the common training and evaluation setting described in Section III.
In the Classical mode, the server-side latent feature is passed directly to the classifier. In the Quantum mode, the same latent feature is additionally processed by the QEP, which produces observable-based quantum statistics, decodes them into the shared latent space, and fuses them with the classical representation. This comparison isolates the contribution of post-aggregation quantum refinement under matched conditions.
We report both predictive performance and internal QEP behavior. Predictive evaluation uses both the standard threshold and the validation-optimized threshold. Internal analysis tracks the mean fusion coefficient $\alpha$, the dispersion of the quantum-enhanced branch through $q$-std, and their training behavior across frontend types.
Because TTN provides the most favorable overall profile in the frontend comparison, we additionally use the TTN-based model to examine the operating range of the QEP under qubit scaling and noise variation. For qubit scaling, the QEP is instantiated with $N_{q}\in\{4,6,8,10,12,14,16\}$, and the resulting test-accuracy distributions are compared with the classical TTN baseline. For noise sensitivity, we fix $N_{q}=8$ and compare the classical TTN baseline with four quantum execution conditions: noiseless, depolarizing, thermal, and mixed noise. Both analyses are reported as distributional summaries of test accuracy over repeated runs.
@@ IV-B Experiment 2: MPC Mode Benchmark
The second experiment benchmarks the communication cost of MPC-secured aggregation under progressively stronger protection settings. The benchmark isolates the communication induced by protected latent aggregation and is not intended as a full accounting of all end-to-end distributed traffic. Rather, it is designed to clarify how protected representation dimension and security setting determine the communication burden of the secure aggregation stage.
We vary the number of clients from $n=1$ to 3030. For each value of $n$, we consider seven scenarios. Scenario 0 is an insecure baseline in which a central server $S$ performs the aggregation and transformation directly. Scenarios 1–3 correspond to passive-security MPC settings with progressively richer protected functionality: Scenario 1 protects only weighted aggregation, Scenario 2 additionally protects normalization, and Scenario 3 further protects the post-aggregation transformation stage. Scenarios 4–6 are the corresponding active-security counterparts. Table I summarizes these settings.
@@ IV-B1 Cost Model and Measurement Protocol
We employ a symbolic cost meter that tracks three types of communication: (i) data sent by clients to computation nodes, (ii) data exchanged between computation nodes, and (iii) data transmitted during output reconstruction.
In all scenarios, we assume $k=64$, corresponding to an encoding in which client inputs fit into 64-bit machine words. For Scenarios 2 and 5, division is metered using $\theta=5$. The benchmark is intended to capture relative scaling trends across MPC scenarios rather than protocol-tight cryptographic cost bounds.
@@ IV-B2 Fairness and Comparability Across Modes
To avoid conflating security modeling with numerical instability, we assume that fixed-point parameters can be chosen so that the accuracy of the MPC computation matches that of the original insecure floating-point computation. As noted in [27], automated tools can assist in fixed-point parameter selection. Under this assumption, the benchmark should be interpreted as a comparison of secure-computation overheads under matched numerical behavior.
@@ V-A Results of Experiment 1: Quantum Enhancement across MPS, TTN, and MERA
Figure 4: Threshold-optimized test performance of the Classical and Quantum modes across MPS, TTN, and MERA. Boxplots report class-wise Precision, Recall, F1-score, and overall Accuracy for the default QEP setting ($N_{q}=16$).
@@
Figure 6: TTN-based QEP: (a) test-accuracy distributions under qubit-count scaling and (b) robustness under representative noise conditions at $N_{q}=8$. The classical TTN baseline is shown for reference.
@@
Figs. 4 and 5 summarize the comparison between Classical and Quantum modes across MPS, TTN, and MERA. Overall, the results show that the effect of quantum enhancement is architecture-dependent rather than uniform across tensor-network frontends.
Under threshold-optimized evaluation, TTN+QEP provides the most balanced overall profile among the tested frontend configurations. Across architectures, the main effect of the QEP is not a uniform increase in Accuracy, but a frontend-dependent redistribution of class-wise Precision, Recall, and F1. Pneumonia Recall remains high across all three frontend types, staying around 0.90 in the present setting.
The internal diagnostics also differ across frontends. The mean fusion coefficient $\alpha$ decreases during training for all architectures, but the resulting operating regimes are distinct. MPS reaches the highest final $q$-std, MERA remains intermediate, and TTN stabilizes at a comparatively lower and more controlled quantum-branch dispersion. These observations suggest that the effect of the QEP depends not only on the quantum branch itself, but also on the latent structure induced by the tensor-network frontend.
@@
Fig. 6 shows additional analyses for the TTN-based QEP. In the qubit-scaling comparison, the lower-qubit settings, especially $N_{q}=4$ and $N_{q}=6$, tend to show less favorable accuracy distributions than the higher-qubit settings. By contrast, from $N_{q}\geq 8$, the distributions appear more stable and remain within a broadly similar range.
This behavior is qualitatively consistent with the scaling consideration introduced in Section II-D. For the present latent dimension $d=64$, the transition to more stable behavior from around $N_{q}=8$ is compatible with the heuristic relation $N_{q}\sim\sqrt{d}$. Although this does not constitute a formal scaling law, it supports the view that the usefulness of the QEP depends on a reasonable match between quantum feature capacity and latent representation size.
@@
The noise analysis at $N_{q}=8$ shows degradation under all tested noisy conditions relative to the noiseless setting. The depolarizing condition remains comparatively close to the noiseless case, whereas the thermal condition shows a wider spread and includes lower-performing runs. The mixed-noise condition also shows a degradation trend while retaining a broadly comparable range.
@@ V-B Results of Experiment 2: MPC Benchmark
Fig. 7(a) and 7(b) summarize the modeled communication overhead for the passive and active MPC scenarios together with the insecure baseline.
@@
The benchmark shows that, under the present abstraction of server-side MPC communication, the dominant communication cost is governed by the bit-width $k$, the protected representation dimension, and the selected protection scenario, rather than by the number of aggregated clients itself. In both figures, $k=64$ is fixed while the protected representation dimension varies from the raw dimension $D_{\mathrm{raw}}=784$ to the compressed latent dimension $d=64$.
Active security preserves the same qualitative scaling trend as passive security while introducing an additional multiplicative overhead. Across both passive and active settings, the dominant communication trend is driven by the representation dimension, indicating that client-side compression directly reduces the communication surface on which secure aggregation is applied.
Because the present benchmark isolates the communication cost of the MPC stage itself, the number of clients does not emerge as the dominant scaling factor in the reported curves. Under the linear secret-sharing model considered here, the main cost trend is governed primarily by protected representation dimension and protection scenario, indicating that client-side compression has direct systems-level value because it reduces the communication surface on which secure aggregation is applied.
@@ VI Discussion
From a clinical perspective, the consistently high Pneumonia Recall across architectures is encouraging. In screening or triage settings, sensitivity to pneumonia-positive cases is often more important than small fluctuations in class-balanced metrics because false negatives carry disproportionate clinical cost.
At the same time, the present results should not be read as establishing TTN+QEP as a universally optimal combination. Rather, they indicate that the utility of post-aggregation quantum refinement is contingent on the latent structure delivered by the frontend and on how that structure matches the effective capacity of the quantum feature map. In this sense, the main lesson is not the dominance of a single architecture, but the importance of co-design across compression, aggregation, and quantum refinement.
@@
A related direction for quantum feature extraction in this framework is to replace or augment the present observable-based QEP with a quantum reservoir readout acting on the aggregated latent input. Quantum reservoir computing is attractive in this context because it exploits the natural nonlinear dynamics of many-body quantum systems while avoiding gradient-based variational optimization of the quantum circuit itself. This is relevant from a practical viewpoint, as variational quantum models can suffer from barren plateau phenomena, whereas quantum kernel methods can also face exponential concentration of kernel values and associated measurement inefficiency under certain conditions [18, 33]. By contrast, reservoir-style quantum processing uses the dynamics of a fixed quantum system as a feature generator and trains only the classical readout, which provides a complementary design point for small- to intermediate-scale hybrid learning. Recent studies have experimentally demonstrated repeated-measurement quantum reservoir computing on superconducting devices and have also reported large-scale analog quantum reservoir learning at the 100-qubit scale, supporting the practical relevance of this direction [37].
The longer-term significance of the architecture lies in future qubit regimes where faithful classical simulation is no longer practical, in which case post-aggregation quantum processing on tensor-network-compressed latent features becomes a more meaningful systems option than direct quantum processing of the original high-dimensional input. This interpretation is also consistent with the qubit-scaling result. In the present setting, improved stability emerges only once the quantum feature capacity becomes reasonably matched to the latent representation size, suggesting that frontend structure and quantum resource scale should be considered jointly rather than independently.
@@ VII Conclusion
In this work, we investigated the role of the QEP within a privacy-aware federated learning framework for medical image classification. The proposed architecture combines client-side tensor-network representation learning, an abstract secure-aggregation model for MPC benchmarking, and a post-aggregation quantum refinement module.
@@
The additional qubit-scaling and noise analyses further clarify the practical behavior of the QEP. For the present 64-dimensional latent input, performance becomes more stable from 88 qubits onward, while noisy conditions at 88 qubits degrade performance relative to the noiseless case in a noise-type-dependent manner. These results do not establish hardware-level benefit, but they help characterize the operating range and robustness of the hybrid quantum branch within the proposed pipeline.
@@
While the present results are simulator-based, an important long-term motivation of the proposed framework is future operation in regimes where classical simulation of the quantum branch is no longer practical. In such settings, post-aggregation quantum processing on tensor-network-compressed latent features may become more meaningful than direct quantum processing of the original high-dimensional input, because compression reduces the representation to a scale that is more compatible with limited quantum hardware while preserving task-relevant structure. From this perspective, the present study should be viewed as a systems-oriented step toward hybrid quantum federated learning under realistic privacy and resource constraints, rather than as a claim of near-term quantum advantage.
Future work will proceed in three directions. First, hardware validation in qubit regimes beyond practical classical simulation will be important for testing whether the proposed TN+QEP interface remains useful beyond the simulator-based setting considered here. Second, evaluation on real medical data will be needed to determine how far tensor-network frontends can compress clinically relevant inputs while preserving task-relevant structure and improving the practicality of downstream secure aggregation. Third, end-to-end implementation of the MPC pipeline, including explicit fixed-point execution, will be necessary to validate the present communication-cost model under realistic deployment conditions, assess the full system behavior of the proposed framework, and, when the post-aggregation quantum stage is externally delegated, examine privacy-preserving delegated quantum-computation mechanisms such as blind quantum computing and related protocols [5, 13, 22].
@@ Acknowledgments
The authors acknowledge helpful conversations with Mark Medum Bundgaad, Miyoji Kakinuki, Akiko Kamigori, Kazufumi Okazaki, Takuya Hasegawa, Tomah Sogabe, Naoki Yamamoto, Eriko Kaminishi, Toshiki Yasuda. Hiroshi Yamauchi acknowledges the support of Yoshi-aki Shimada, Yosuke Komiyama and Ryuji Wakikawa. This work was partly supported by NEDO Challenge Quantum Computing “Solve Social Issues!”.